OpenStudy (anonymous):

Solve -15(4+3x)-8x=26-6(x-9)

OpenStudy (ghazi):

expand the equation

OpenStudy (ghazi):

-15*4+(-15)*3x-8x= 26- 6x+54...solve for x

OpenStudy (anonymous):

-15+-15=30 is that right

OpenStudy (anonymous):

30*4=120

OpenStudy (ghazi):

-60-45x-8x= 80-6x.....-140= -59x....just solve for x now

OpenStudy (anonymous):

-60-45x=-15x-8x=-7 80-7=87x 87-6x=93

OpenStudy (anonymous):

-140/-59x=2.37

OpenStudy (anonymous):

its that wrong

OpenStudy (ghazi):

yes... the answer is x= 140/59